OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) <--> 2x − 6

OpenStudy (anonymous):

both are equal...

OpenStudy (campbell_st):

the left hand side uses the distributive power over multiplication 2(x - 3) = 2*x - 2*3 = 2x -3 so the left hand side equals the right hand side.